# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ Abstract
Interactive agents capable of understanding and executing instructions in the physical world have long been a central goal in AI research. The Minecraft Collaborative Building Task (MCBT) provides one such setting to work towards this goal (Narayan-Chen, Jayannavar, and Hockenmaier, 2019). It is a two-player game in which an Architect (A) instructs a Builder (B) to construct a target structure in a simulated Blocks World Environment. We focus on the challenging Builder Action Prediction (BAP) subtask of predicting correct action sequences in a given multimodal game context with limited training data (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). We take a closer look at evaluation and data for the BAP task, discovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it. It comprises of: (1) an enhanced evaluation benchmark that includes a cleaner test set and fairer, more insightful metrics, and (2) additional synthetic training data generated from novel Minecraft dialogue and target structure simulators emulating the MCBT. We show that the synthetic data can be used to train more performant and robust neural models even with relatively simple training methods. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. Although modeling is not the primary focus of this work, we also illustrate the impact of our data and training methodologies on a simple LLM- and transformer-based model, thus validating the robustness of our approach, and setting the stage for more advanced architectures and LLMs going forward.
@@ 1 Introduction
There has long been interest in developing interactive agents that can communicate with humans about and operate within the physical world (e.g., Winograd (1971)). The goal for these agents is not only to engage in rich natural language dialogue, but also to ground it to physical objects, and execute instructions in the real world. We work toward this goal using Minecraft (https://minecraft.net/), a popular multiplayer game where players control avatars to navigate a 3D world and build structures from block-like materials. Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform. Our previous work—the Minecraft Collaborative Building Task (MCBT) and the corresponding Minecraft Dialogue Corpus (MDC) Narayan-Chen, Jayannavar, and Hockenmaier (2019)—was among the first to use it to study grounded task-oriented dialogue. Jernite et al. (2019) is another such example. More recent works, such as Mohanty et al. (2024), have been directly inspired by the MCBT. Additionally, there now exists a well-developed set of platform and tools that support AI experimentation within Minecraft, including Malmo Johnson et al. (2016), Craftassist Gray et al. (2019), TaskWorldMod Ogawa et al. (2020), MC-Saar-Instruct Köhn et al. (2020), IGLU GridWorld Zholus et al. (2022), and Narayan-Chen, Jayannavar, and Hockenmaier (2019). The latter builds on the Malmo platform.
@@
The MCBT situates a dialogue task—designed to be asynchronous, asymmetric, and with minimal language constraints—in a simulated 3D Minecraft environment where participants, with constantly shifting perspectives, must refer to and modify a dynamic world. In this task, two players collaborate: an Architect (A) instructs a Builder (B) to construct a target structure using multi-colored blocks (Figure 1). The MDC comprises 509 human-human game logs for this task. Narayan-Chen, Jayannavar, and Hockenmaier (2019) focus on generating Architect utterances, and Jayannavar, Narayan-Chen, and Hockenmaier (2020) focus on end-to-end neural models for building an automated Builder agent. The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game. Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via data augmentation techniques) yield promising initial results. We recap relevant prior work on the Minecraft Builder in Section 2. Building on this foundation, we next take a closer look at evaluation and training data for the BAP task, uncovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it.
Given the inherent complexity of the task, we find that evaluation is nuanced and requires closer examination. We identify key challenges that hinder fair and insightful evaluation, then systematically address them. To this end, we introduce a new, cleaner v2 test set to replace the original legacy test set for evaluating BAP models (Section 3). We also refine the F1 metric, proposing a fairer variant that accounts for nuances in the data. The current evaluation, which relies solely on an aggregated and opaque F1 metric, lacks detailed insights into model behavior. To enhance insightfulness, we introduce additional metrics that measure specific model capabilities. Together, these contributions form the updated BAP v2 evaluation benchmark (Section 4).
Our work on evaluation highlights spatial reasoning as a key bottleneck in model performance, alongside the challenge of limited training data (even with the aforementioned data augmentation). Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5). This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with straightforward training methods (Section 6). We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance. Further, a straightforward Curriculum Learning method provides an additional boost. Together, the synthetic and augmented BAP data form the BAP v2 training set. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. We provide a detailed analysis of model performance of our best performing model relative to the baseline in Section 7.
Although modeling is not the primary focus of this work, solely to test the robustness of our synthetic data and training methodologies and show their applicability to relatively contemporary model architectures, we experiment with very simple LLM (BERT)- and transformer-based variants of our neural models. The simplicity also allows us to assess whether/to what extent simpler LLMs like BERT and vanilla Transformers alone can already effectively handle the BAP task. Our results show that the synthetic data and training methodologies improve these models as well, validating the robustness of our approach (Section 6.3). Future work can build on these insights to explore more sophisticated LLMs and architectures while leveraging the BAP v2 framework.
Section 8 provides a detailed discussion on four topics – an additional key challenge in BAP evaluation, the importance of evaluation on the synthetic data, relevant concurrent work, and broader implications of our work that go beyond BAP. The latter further highlights the potential impact of our work and directions for future research. Related work is covered in Section 9. We conclude and discuss future work in Section 10. We plan to release code and data involved in this work.
@@ 2 Prior Work on the Minecraft Builder
This section provides a high-level overview of the dataset, task, and baseline model, summarizing our prior work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020). We also introduce some necessary formalisms and figures for better exposition.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
Figure 2: A sample sequence of human-human game states. The game starts with an empty board and an initial A instruction (a), which B executes in the first action sequence (b) by placing a single block. In (c), B begins to execute the next A instruction given in (b). However, A interrupts B in (c), leading to two distinct B action sequences: (b)–(c) (single block placement), and (c)–(h) (multiple placements and removals).
@@ 2.1 The Minecraft Collaborative Building Task and Dialogue Corpus
2.1.1 Features and Challenges
Since target structures can be fairly complex, Architects typically give step-by-step instructions ("now a yellow 6") for different parts of the target. Builders should execute these instructions, but may also ask questions ("Where does the 6 start") which the Architects has to answer ("behind the 7 from your perspective"). Architects may also need to identify and correct mistakes the Builder may have made ("too much overlap unfortunately"). A can move around freely, but remains invisible to B and views the structure from behind B when giving instructions. As a result, A instructions frequently include spatial relations, both between pairs of blocks or substructures (“put … on top of…"), and relative to B’s current position and perspective (“left", “right"). Humans also often use higher-level descriptions to refer to complex (sub)shapes (e.g. “staircase", “v"). Due to the asynchronous nature of the dialogue, the Architect often talks while the Builder is placing blocks (“so it will look like a v"), leading to an apparent interruption of the Builder’s action sequences. Figure 2 shows an example from the MDC that highlights some of these challenges.
Floating blocks
Blocks do not need to be placed on the ground if their cell is adjacent to a supporting block. If that supporting block is later removed, the remaining block (and any structure supported by it) will “float” in place. Thus, placing floating blocks needs the placement and subsequent removal of such placeholder supporting blocks. Instructions for floating structures vary greatly, ranging from step-by-step instructions involving temporary supporting blocks to single-shot descriptions such as, simply, “place a floating yellow block”.
@@ 2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
We define the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart. We note that compared to BAP, Shi, Feng, and Lipani (2022) propose a slightly different building task involving predicting a single action instead of a sequence of actions. Thus, their setup is not directly comparable to ours.
@@ 2.2 The Builder Action Prediction (BAP) Task
The Builder Action Prediction (BAP) task is defined as the task of predicting the actions (block placements and/or removals) that a human Builder performed at a particular point in a human-human game. Games start with an empty board, and consist of an alternating sequence of dialogues $D^{t}$ (which may each consist of a single utterance, or a back-and-forth exchange between the two players) and Builder actions $A^{t}$ that result in an updated built structure $S^{t}$. Each action sequence $A^{t}$ yields a BAP item $(\mathcal{H}^{t-1},S^{t-1},D^{t},A^{t},S^{t})$ that consists of the game history $\mathcal{H}^{t-1}$ that culminated in the previous built structure $S^{t-1}$, the new dialogue $D^{t}$, and the new action sequence $A^{t}$ that resulted in the new structure $S^{t}$. Given $\mathcal{H}^{t-1}$, $S^{t-1},$ and $D^{t}$, BAP models $M$ should predict an action sequence $A^{t}_{M}$ that yields $S^{t}$.
Dataset statistics
The training, test and development splits of the MDC contain 3709, 1616, and 1331 items respectively, and the average sequence length of an action sequence (in the development set) is 4.3 (with a std. deviation of 4.5). Target structures in the test data do not appear in the training or development data.
Example
@@
To illustrate this task, Figure 3 shows how the game snippet shown in Figure 2 corresponds to three different items for the BAP task. The leftmost columns in each figure display the (truncated) game history ($S^{t-2}$, $D^{t-1}$, $S^{t-1}$, and $D^{t}$), while the rightmost column shows the resulting built structure $S^{t}$. Columns are read top to bottom and left to right. Note that in this figure (and similar ones in this work), we use $S_{H}$ instead of $S$ for clarity ($H$ indicates that the structure was built by a human, not a model). To highlight the effect of the final action sequence, previously placed blocks are shown in lighter colors, and each structure is rotated based on B’s gaze (only the yaw angle) at the end of the dialog snippet below it. B’s exact position and pitch angle are omitted for clarity. Any temporary supporting blocks are not shown as well. The game starts with an empty board $S^{0}$, and an initial dialogue between the two players $D^{1}$, and the first BAP item consists of predicting that first block placement $A^{1}$ (and consequent structure $S^{1}$) (Figure 3(a)). Then, another instruction follows, which the Builder executes by placing another block, resulting in the next structure to be predicted (Figure 3(b)), and the third instruction yields a complex action sequence that results in the desired v-shape structure (Figure 3(c)).
((a)) Item 1
((b)) Item 2
((c)) Item 3
Figure 3: The snippet in Figure 2 corresponds to three BAP task items as shown.

### v3

@@ Abstract
Developing interactive agents that can understand language, perceive their surroundings, and act within the physical world is a long-standing goal of AI research. The Minecraft Collaborative Building Task (MCBT) (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), a two-player game in which an Architect (A) instructs a Builder (B) to construct a target structure in a simulated 3D Blocks World environment, offers a rich platform to work towards this goal.
In this work, we focus on the Builder Action Prediction (BAP) subtask: predicting B’s actions in a multimodal game context (Jayannavar, Narayan-Chen, and Hockenmaier, 2020) – a challenging testbed for grounded instruction following, with limited training data. We holistically re-examine this task and introduce BAP v2 to address key challenges in evaluation, training data, and modeling. Specifically, we define an enhanced evaluation benchmark, featuring a cleaner test set and fairer, more insightful metrics that also reveal spatial reasoning as the primary performance bottleneck. To address data scarcity and to teach models basic spatial skills, we generate different types of synthetic MCBT data. We observe that current, LLM-based SOTA models trained on the human BAP dialogues fail on these simpler, synthetic BAP ones, but show that training models on this synthetic data improves their performance across the board. We also introduce a new SOTA model, Llama-CRAFTS, which leverages richer input representations, and achieves an F1 score of 53.0 on the BAP v2 task and strong performance on the synthetic data. While this result marks a notable 6 points improvement over previous work, it also underscores the task’s remaining difficulty, establishing BAP v2 as a fertile ground for future research, and providing a useful measure of the spatial capabilities of current text-only LLMs in such embodied tasks.
@@ 1 Introduction
Developing interactive agents that can understand language, perceive their surroundings, and act within the physical world is a long-standing goal of AI research (e.g., Winograd (1971)). We work toward this goal in Minecraft (https://minecraft.net/), a popular 3D game that has become a prominent platform for AI experimentation. Our previous work introduced the Minecraft Collaborative Building Task (MCBT) and the corresponding Minecraft Dialogue Corpus (MDC) Narayan-Chen, Jayannavar, and Hockenmaier (2019), one of the first efforts to use this environment to study grounded, task-oriented dialogue.
@@
In the MCBT, two players collaborate to construct a target structure from multi-colored blocks (Figure 1): an Architect (A) provides instructions to a Builder (B). The task is situated in a dynamic 3D world that participants must refer to and modify from constantly shifting perspectives. The dialogue is asynchronous, asymmetric, and has minimal linguistic constraints. The MDC contains 509 human-human game logs from this task. While Narayan-Chen, Jayannavar, and Hockenmaier (2019) focused on generating Architect utterances, Jayannavar, Narayan-Chen, and Hockenmaier (2020) introduced end-to-end models for an automated Builder. The latter work defined the challenging Builder Action Prediction (BAP) task: predicting the sequence of actions (block placements and/or removals) that a human Builder performed at a given point in a human-human game. We review prior work on the Minecraft Builder and related recent works in Section 2.
The BAP task remains a relevant and challenging testbed for grounded instruction following, presenting a unique combination of difficulties. Unlike navigation or object-centric benchmarks such as VLN (Anderson et al., 2018), CVDN (Thomason et al., 2019), ALFRED (Shridhar et al., 2020), DialFRED (Gao et al., 2022), and TEACh (Padmakumar et al., 2022), where agents primarily interact with existing objects, the MCBT requires grounding instructions that refer to objects and structures which do not yet exist, demanding greater abstraction and planning. The dialogue is also more complex than in un-embodied benchmarks like MultiWoz (Budzianowski et al., 2018) or other Minecraft datasets like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a); it features free-form, asynchronous communication within a world that is dynamically modified. Finally, the task poses a significant planning problem due to its large action space—over 7,600 possible actions compared to just six in typical VLN work—and its reliance on complex, perspective-dependent spatial relations (e.g., “to your left"), which are more difficult to resolve than the absolute cardinal directions used in IGLU. This combination of challenges makes BAP a compelling setting for research.
However, to enable future research to make more efficient and meaningful progress, this paper holistically re-examines the BAP task framework. In particular, we identify and address key challenges in evaluation, training data, and modeling, and propose BAP v2, an upgraded version of the task: [^1] All code and data developed for this research will be released at https://github.com/prashant-jayan21/bap-v2
Evaluation (Sections 3 and 4)
The complexity of the BAP task makes evaluation nuanced. We systematically address challenges that hinder fair and insightful assessment by introducing a cleaner v2 test set (Section 3), propose a fairer F1 metric and introduce new auxiliary metrics to measure specific model capabilities (Section 4). Together, these contributions establish the BAP v2 evaluation benchmark and help us move beyond the original, opaque F1 score, providing deeper insights. Adopting this new benchmark also reveals that spatial reasoning is the key performance bottleneck in the BAP task.
Training Data (Section 5)
To overcome the dual challenges of limited training data for the BAP task—stemming from the small size of the MDC (509 game logs)—and the key performance bottleneck of spatial reasoning, we create a diverse set of synthetic training data. This data is generated using novel Minecraft dialogue and target structure simulators that were designed emulate the MCBT. While simpler than the human MDC data, the synthetic dialogues are rich in the spatial relations and referring expressions crucial for this task. The combination of this new data with the original corpus forms the BAP v2 training set. Our approach builds on established work showing the effectiveness of synthetic data for various tasks, from vision-and-language navigation to dialogue systems (Kamath et al., 2023; Ku et al., 2020; Wang et al., 2023; Kang et al., 2023; Das et al., 2017; Kim et al., 2022; Bao et al., 2023; Zhan et al., 2023, 2024; Tang et al., 2024), although synthesizing complete, task-oriented embodied dialogues remains a relatively underexplored area (Padmakumar et al., 2023), especially since the MCBT is significantly more complex than other Blocks World settings (Bisk, Yuret, and Marcu, 2016; Dan, Han, and Roth, 2021).
Modeling and Analysis (Sections 6 and 7)
To demonstrate the utility of our synthetic data in addressing the challenge of limited original BAP data, we show that it produces more performant models, for both LLM and non-LLM architectures, even with straightforward training methods (Section 6). The data also reveals that both LLM and non-LLM models trained solely on the complex BAP human dialogues lack robustness, counterintuitively underperforming on our simpler synthetic data. As the final component of the BAP v2 framework, we extend prior work on LLM-based models for the BAP task (Chaturvedi, Thompson, and Asher, 2024) by introducing richer input text representations that capture essential game context, and we find that these enhanced representations and the synthetic data are both necessary and mutually reinforcing. Our best model, Llama-CRAFTS (Context Rich And Fine-Tuned On Synthetic Data), establishes a new state of the art on the BAP task as well as achieves strong performance on the synthetic data. But while its F1 score of 53.0 on BAP represents a 6 point improvement over the previous SOTA (Chaturvedi, Thompson, and Asher, 2024) (under the new v2 evaluation), it is nevertheless clear that there is much room for further improvement. A detailed analysis of model output (Section 7) illustrates how Llama-CRAFTS improves over the previous SOTA (especially in spatial reasoning), where it falls short, what this implies for current text-only LLMs’ spatial capabilities in such embodied tasks, and indicates both strengths and limitations of our new evaluation metrics.
While spatial reasoning in language models has been explored in text and across other modalities (Mirzaee and Kordjamshidi, 2022; Shi, Zhang, and Lipani, 2022; Rizvi, Zhu, and Gurevych, 2024; Gopinathan et al., 2024; Li et al., 2024a; Bisk et al., 2018; Du et al., 2024; Kamath, Hessel, and Chang, 2023; Wang et al., 2024), our work addresses a specific gap in the literature: exploring the capabilities of text-only LLMs in a 3D, embodied, collaborative dialogue setting. Prior work has shown LLMs struggle with even simple 2D grid navigation and path-planning in text (Yamada et al., 2024; Aghzal, Plaku, and Yao, 2024), and has explored spatial reasoning in games like chess (Feng et al., 2023a; Zhang et al., 2025b) or puzzles (Chollet et al., 2025), but no existing research combines this with the challenges of 3D grids in text, and embodied, collaborative dialogue, thus motivating our focus.
Part 4: Looking Ahead (Sections 8 and 9)
Finally, we discuss future research directions (Section 8) We outline the key next challenges for BAP that can be tackled by building on the BAP v2 framework, and explore the broader implications of our work beyond BAP, including way in which ideas from our synthetic data and simulators hold the potential for furthering other related works and tasks (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Mohanty et al., 2024; Bonn et al., 2020; Thompson, Hunter, and Asher, 2024; Bonial et al., 2021; Madge et al., 2025). We conclude in Section 9.
@@ 2 Prior Work on the Minecraft Builder
Sections 2.1,2.2,2.3,2.5 and 2.6 provide a high-level overview of the Minecraft Collaborative Building Task and Dialogue corpus, the BAP task and dataset, evaluation metrics, and baseline model, summarizing our prior work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Section 2.4 formalizes the BAP task. Section 2.7 discusses recent related work around more contemporary LLM-based models for BAP (Chaturvedi, Thompson, and Asher, 2024; Kranti, Hakimov, and Schlangen, 2024), and Section 2.8 discusses related work around the task and dataset itself, explaining why the BAP task continues to be a relevant and challenging testbed for grounded instruction following.
A snippet from the MDC corpus
Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d)
Refer to caption Refer to caption Refer to caption Refer to caption
(e) (f) (g) (h)
Figure 2: A sample sequence of human-human game states from the MDC corpus. The game starts with an empty board and an initial A instruction (a), which B executes in the first action sequence (b) by placing a single block. In (c), B begins to execute the next A instruction given in (b). But A interrupts B in (c), leading to a new B action sequence in (c)–(h) (multiple placements and removals).
@@ 2.1 The Minecraft Collaborative Building Task and Dialogue Corpus
2.1.1 Features and Challenges of the MDC
Since target structures can be fairly complex, Architects typically give step-by-step instructions (“now a yellow 6") for different parts of the target. Builders should execute these instructions, but may also ask questions (“Where does the 6 start") which the Architects has to answer (“behind the 7 from your perspective"). Architects may also need to identify and correct mistakes the Builder may have made (“too much overlap unfortunately"). A can move around freely, but remains invisible to B and views the structure from behind B when giving instructions. As a result, A instructions frequently include spatial relations, both between pairs of blocks or substructures (“put … on top of…"), and relative to B’s current position and perspective (“left", “right"). Humans also often use higher-level descriptions to refer to complex (sub)shapes (e.g. “staircase", “v"). Due to the asynchronous nature of the dialogue, the Architect often talks while the Builder is placing blocks (“so it will look like a v"), leading to an apparent interruption of the Builder’s action sequences. Finally, Minecraft, blocks do not need to be placed on the ground if their cell is adjacent to a supporting block. If that supporting block is later removed, the remaining block (and any structure supported by it) will “float” in place. Thus, placing floating blocks needs the placement and subsequent removal of such placeholder supporting blocks. Instructions for floating structures vary greatly, ranging from step-by-step instructions involving temporary supporting blocks to single-shot descriptions such as, simply, “place a floating yellow block". Figure 2 shows an example from the MDC that highlights some of these features and challenges.
@@ 2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
Although the MDC should ultimately contribute to the creation of fully interactive agents that can complete an entire game, we have found it helpful, and even necessary, to start with simpler tasks. Specifically, we have defined the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart.
Three BAP task items corresponding to the snippet in Figure 2
Item 1 corresponding to (a)–(b) in Fig. 2
Item 2 corresponding to (b)–(c) in Fig. 2
Item 3 corresponding to (c)–(h) in Fig. 2
Figure 3: The snippet in Figure 2 corresponds to three BAP task items shown here.
@@ 2.2 The Builder Action Prediction (BAP) Task
The Builder Action Prediction (BAP) task is defined as the task of predicting the actions (block placements and/or removals) that a human Builder performed at a particular point in a human-human game. Games start with an empty board, and consist of an alternating sequence of dialogues $D^{t}$ (which may each consist of a single utterance, or a back-and-forth exchange between the two players) and Builder actions $A^{t}$ that result in an updated built structure $S^{t}$. Each action sequence $A^{t}$ yields a BAP item $(\mathcal{H}^{t-1},S^{t-1},B^{t},D^{t},A^{t},S^{t})$ that consists of the game history $\mathcal{H}^{t-1}$ that culminated in the previous built structure $S^{t-1}$, the Builder’s position and orientation, $B^{t}$, the new dialogue $D^{t}$, and the new action sequence $A^{t}$ that resulted in the new structure $S^{t}$. Given $\mathcal{H}^{t-1}$, $S^{t-1},$ and $D^{t}$, BAP models $M$ should predict an action sequence $A^{t}_{M}$ that yields $S^{t}$.
@@
Figure 3 illustrates how the game snippet shown in Figure 2 corresponds to three different items for the BAP task. The game starts with an empty board $S^{0}$, and an initial dialogue between the two players $D^{1}$, and the first BAP item consists of predicting that first block placement $A^{1}$ (and consequent structure $S^{1}$) (Figure 3, Item 1). Then, another instruction follows, which the Builder executes by placing another block, resulting in the next structure to be predicted (Figure 3, Item 2), and the third instruction yields a complex action sequence that results in the desired v-shape structure (Figure 3, Item 3).
How to interpret BAP item figures in this paper
The leftmost columns in each BAP item figure (e.g. Figure 3) display the (truncated) game history ($S^{t-2}$, $D^{t-1}$, $S^{t-1}$, and $D^{t}$), while the rightmost column shows the resulting built structure $S^{t}$. Columns are read top to bottom and left to right. We use $S_{H}$ instead of $S$ to indicate that the structure was built by a human ($H$), not a model). To highlight the effect of the final action sequence, previously placed blocks are shown in lighter colors. Structures like $S^{t-1}$ are rotated to align with the B’s POV $B^{t}$ (corresponding to the next turn and dialogue segment $D^{t}$ below it). For B’s POV, we only use the yaw angle, and B’s exact position and pitch angle are omitted for clarity. Any temporary supporting blocks are not shown as well.
2.3 The original BAP dataset
The training, test and development splits of the MDC contain 3709, 1616, and 1331 BAP items respectively, and the average sequence length of an action sequence (in the development set) is 4.3 (with a standard deviation of 4.5). Target structures in the test data do not appear in the training or development data. Since the small size of the training set is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three simple data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B). Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for the GRU-based baseline model (Section 2.6). Henceforth, we will refer to this combination of the augmented training data and the original test and development data as the original BAP data.